OpenStudy (anonymous):

Can someone explain me how to do this GPA % of Juniors 0.00 – 1.99 16% 2.00 – 2.49 24% 2.50 – 2.99 28% 3.00 – 3.49 22% 3.50 – 4.00 10% Referring to the data above, find the probability that a randomly selected junior's GPA is in the 2.50 - 3.49 range, given that it is greater than 1.99. the answer should be .676

OpenStudy (anonymous):

if assume that X is the height; P(2.50<X<3.49 | X>1.99) = P( 2.50<X<3.49 intersection X>1.99)/P(X>1.99) the intersection is :-

OpenStudy (anonymous):



OpenStudy (anonymous):

P( 2.50<X<3.49 intersection X>1.99)= > P( 2.50<X<3.49 )

OpenStudy (anonymous):

this is the new one P( 2.50<X<3.49 )/ P(X>1.99) = (50/100)/(84/100) = 0.5952 this is the answer